OpenStudy (anonymous):

another limits question, please help!! will be typed in equation editor in a few mins.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\lim_{x \rightarrow 64} [x-64]\div \sqrt[3]{x}-4\]

OpenStudy (anonymous):

is that 3rd root of x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

see here we rationalize the denom to get rid of that 3rd root

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

or actually, we can use hopitals rule...

OpenStudy (anonymous):

i don't know what that is :/

OpenStudy (anonymous):

when u have a limit of the form 0/0 by plugging in the limiting value, we can take the derivative of the top and derivative of the bottom and re-write the fraction and do the limit again

OpenStudy (anonymous):

in this case: = lim x--> 64 (1 / (1/3)x^(-2/3)) and by plugging in 64 (bcuz we can now) we get: = 48

OpenStudy (anonymous):

oh, i havent learned it that way, so i don't know if it would be accepted on my test.. do you mind showing me how to do it normally?

OpenStudy (anonymous):

ok well the quick way is to simplify x-64 as a difference of cubes: \[\left( x-64 \right)=(\sqrt[3]{x}-4)(x^{2/3}+4\sqrt[3]{x}+16)\] and then plug in 64

OpenStudy (anonymous):

since the denominator will cancel out leaving the trinomial...